OpenStudy (anonymous):

f(x) = sin^4(x+pi/6) + cos^4 (x+pi/6) is a) not a periodic fn b) a periodic fn with period 2pi c) a periodic fn with period pi d) a periodic fn with period pi/2 2. For the hyperbola 9x^2-16y^2-18x+32y-151=0 a) one of the directrix is x =21/5 b) length of latus rectum = 9/2 c) focii are (6,1) & (-4,1) d) ecentricity is 5/4

OpenStudy (anonymous):

help........

OpenStudy (anonymous):

any1 help me........

OpenStudy (anonymous):

ishaan cann u

OpenStudy (anonymous):

sebentar yahhhh tak solve dulu sabar ya jang

OpenStudy (anonymous):

noras please type in english

OpenStudy (anonymous):

waduh soryy je aku ndak iso boso londho je..iki boso jowo lho..indonesia made in punya lah

OpenStudy (anonymous):

please type in english noras

OpenStudy (nilankshi):

hmmmmmmm

OpenStudy (nikvist):

1) \[f(x)=\sin^4(x+\pi/6)+\cos^4(x+\pi/6)=\]\[=\sin^4(x+\pi/6)+2\sin^2(x+\pi/6)\cos^2(x+\pi/6)+\cos^4(x+\pi/6)-\]\[-2\sin^2(x+\pi/6)\cos^2(x+\pi/6)=(\sin^2(x+\pi/6)+\cos^2(x+\pi/6))^2-\]\[-2\sin^2(x+\pi/6)\cos^2(x+\pi/6)=1-\frac{1}{2}\sin^2(2x+\pi/3)=\] \[=1-\frac{1}{2}\frac{1-\cos{(4x+2\pi/3)}}{2}=\frac{3}{4}+\frac{1}{4} \cos{(4x+2\pi/3)}\] period is \[T=2\pi/4=\pi/2\]

OpenStudy (nikvist):

2) \[9x^2-16y^2-18x+32y-151=0\] \[9x^2-18x+9-16y^2+32y-16-144=0\] \[9(x-1)^2-16(y-1)^2=144\] \[\frac{(x-1)^2}{4^2}-\frac{(y-1)^2}{3^2}=1\]

OpenStudy (anonymous):

u on facebook nikvist

OpenStudy (nikvist):

I haven't got account on facebook

OpenStudy (anonymous):

gmail ?

OpenStudy (nikvist):

yes, nikvist83@gmail.com

OpenStudy (anonymous):

please wait for one minute